It is very useful to learn to estimate the answers to science and engineering problems quickly, without doing detailed calculations. As an example, consider a car moving at 3 MPH that hits a brick wall. A "good" bumper should bring the car to a stop in about 2 inches. Estimate the weight of a typical car, and figure this out, making the simple assumption that the force (and thus the acceleration) is constant. The force is about? Explain you answer please
Answer #1

Initial velocity, u = 3mph = 1.34m/s
Final Velocity, v = 0
Distance before Stop, s = 2 in = 0.051 m

v^2 = u^2 + 2*a*s
0 = 1.34^2 + 2*a*0.051
a = - 17.6 m/s^2
(-ve sign means that the acceleratio is in the opposite direction of motion.)

Accelereation , a = 17.6 m/s^2

Force = Mass * Acceleration
Let us imagine a typical car weight , m = 1400 Kg

Force = 1400 * 17.6 N
Force = 24640 N
Force on the Car, F = 24640 N